Solve by graphing if necessary estimate answers to the nearest half unit.
x - 2y = - 6
y = - 3/2 x - 1
: Solve by graphing if necessary estimate answers to the nearest half unit.
x - 2y = - 6
y = - 3/2 x - 1

Answer by jojo14344(1022)

Both are Line Equations---system(x-2y=-6,y=-3x/2-1)
It seems you want them to show in the graph.
In the 1st eqn, as we follow Slope-Intercept Form,
2y=x+6 --->cross(2)y/cross(2)=(x+6)/2
y=(1/2)x+(6/2)
Reduced -->y=(1/2)x+3, Slope=1/2, y-intercept=3
f(y)=0:
(1/2)x=-3 ---->cross(1/2)x/cross(1/2)=-3/(1/2)
x=-3(2/1)=-6
.
2nd eqn:
y=-3x/2-1 ---> already in the format: Slope=-3/2, y-intercept=-1
F(y)=0:
(-3/2)x=1 ----> cross(-3/2)x/cross(-3/2)=1/(-3/2)
x=-2/3
As we see below,
drawing(300,300,-7,7,-4,7,grid(1),graph(300,300,-7,7,-4,7,(1/2)x+3,(-3/2)x-1),circle(-6,0,.15),circle(0,3,.15),circle(-2/3,0,.15),circle(0,-1,.15)) ---> RED Line >>>> x-2y=-6; GREEN Line>>>>> y=(-3/2)x-1
Thank you,
Jojo